Block B has a mass of 3.60kg and is moving to the left at a speed of 3.80m/s. Block A has a mass of 7.20kg and is moving to the right. The two blocks undergo a perfectly inelastic collision. What should be the velocity of Block A in order to have the two blocks remain at rest after the collision?
Answer #1


m1 = 3.6 kg, u1 = -3.8 m/s, m2 = 7.2 kg, v = 0

for inelastic collision

from conservaiton of momentum

m1u1 +m2u2 = (m1+m2)v

3.6*-3.8 +7.2*u2 = (3.6+7.2)*0

u2 = 1.9 m/s